# Physics-informed Active Polarimetric 3D Imaging for Specular Surfaces

### v1

@@ 3 Experiments and Discussion
Conclusion and discussion: The proposed method demonstrates strong generalization and robustness under single-shot acquisition while keeping the high accuracy. It achieves geometrically consistent full-field reconstruction even for surfaces with high spatial frequency structures and large curvature, where conventional single-shot methods fail to provide reliable measurement. Compared with previous analytical approaches, our method offers improved stability and faster inference in practical scenarios.
@@
Despite our encouraging first results, several directions remain for further improvement. Although the network was trained on synthetic data simulated from the digital twin, extending the training data to incorporate a wider range of real sensor characteristics would further enhance generalization. Future work may explore improved sensor-level modeling and hybrid training strategies that combine synthetic and real measurements. In addition, the current framework is specifically designed for specular surfaces. Polarimetric 3D imaging is intrinsically influenced by material properties, and different materials may exhibit distinct polarization responses governed by different reflection models. Extending the proposed approach to handle a broader class of materials, including mixed or spatially varying reflectance, remains an important direction for future research. We hope the proposed method will contribute to more practical and deployment-efficient 3D imaging solutions for complex specular surfaces in dynamic environments.
Disclosures. The authors declare no conflicts of interest.
Data availability. Data underlying the results presented in this Letter are available upon reasonable request.

### v2

@@ IV Conclusion and discussion:
The proposed method demonstrates generalization and robustness under single-shot acquisition while keeping the high accuracy. It achieves geometrically consistent full-field reconstruction even for surfaces with high spatial frequency structures and large curvature, where conventional single-shot methods fail to provide reliable measurement. Compared with previous analytical approaches, our method offers improved stability and faster inference in practical scenarios.
Despite these encouraging results, several directions remain for further improvement. Acquiring reliable ground truth for real specular surfaces is inherently challenging, therefore, the network is trained on synthetic data generated from a digital twin. As a result, the performance on real measurements is slightly lower than that on synthetic data. This discrepancy can be attributed to real polarization camera factors that are not fully simulated by the current rendering model. Extending the training data to incorporate more realistic sensor characteristics would further enhance generalization. Future work may explore improved sensor-level modeling as well as hybrid training strategies that combine synthetic and real data.
@@
In addition, the current framework is specifically designed for specular surfaces. Polarimetric 3D imaging is intrinsically influenced by material properties, and different materials may exhibit distinct polarization responses governed by different reflection models. Extending the proposed approach to handle a broader class of materials, including mixed or spatially varying reflectance, remains an important direction for future research. We hope the proposed method will contribute to more practical and deployment-efficient 3D imaging solutions for complex specular surfaces in dynamic environments.
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request..
Conflict of Interest Statement
The authors have no conflicts to disclose.